Circular No. 6152

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

ALEXIS J1139–685 D. Roussel-Dupré, J. J. Bloch, B. C. Edwards, T. E. Pfafman, W. C.

Priedhorsky, S. Ryan, and B. W. Smith, Astrophysics and Radiation Measurements group, Los Alamos National Laboratory; and O. H. W. Siegmund, S. Cully, T. Rodriguez-Bell, J. Vallerga, and J. Warren, University of California at Berkeley, report: “Around Mar. 17.5 UT, the ALEXIS extreme-ultraviolet satellite observed the beginning of a short-lived transient that peaked in flux on Mar. 18.0. By Mar. 18.5, the transient was no longer observable by ALEXIS. The location of the transient is α = 11^{h}39^{m}, δ = −68^{o}30′ (equinox 2000.0; accuracy ± 0^{o}.5). The transient was observable in the narrow 70-eV channel (17.6 nm) but not the narrow 93-eV channel (13.3 nm). The approximate observed count rate, integrated over the burst duration, implies a continuum flux of 0.03 ± 0.008 photon cm^{−2} s^{−1} A^{−1} or line emission of 0.44 photon cm^{−2} s^{−1} at the 70-eV band center. A quick search of various stellar catalogues did not produce any likely candidates within the error box.”

X-RAY NOVA 1994 IN SCORPIUS C. A. Wilson, B. A. Harmon, S. N. Zhang, W. S. Paciesas, and G. J.

Fishman report for the Compton Observatory BATSE Team: “The xray/radio transient GRO J1655–40 is undergoing a major outburst in hard x-rays above 20 keV. Preliminary fluxes for the past several days in the band 20–100 keV are: Mar. 17, 240 ± 20 mCrab; 18, 470 ± 30; 19, 500 ± 30; 20, 480 ± 40; 21, 1200 ± 50. The source has been detectable since mid-February (IAUC 6147), but generally at lower intensity (50–300 mCrab). The spectrum extends to at least 200 keV and is well fit by a power law with spectral index varying between −2.5 and −3.1. The spectrum tends to soften as the intensity increases.”

COMET 29P/SCHWASSMANN-WACHMANN 1 This comet is evidently undergoing another outburst in brightness, as

indicated by the following total visual magnitude estimates: Feb. 10.74 UT, 14.0 (A. Nakamura, Kuma, Ehime, Japan, 0.60-m reflector + unfiltered CCD); 19.65, 14.1 (Nakamura); 24.59, 13.6 (Nakamura); Mar. 8.65, 13.7 (Nakamura); 23.17, 13.3 (A. Hale, Cloudcroft, NM, 0.41-m reflector).

1995 March 23 Daniel W. E. Green